Three resistor are connected in parallel across 100-volt supply. One resistor absorb 1600 watts and the second 1200 watts. The total energy used by the whole circuit in 10 minutes is 276x10^4 joules. Find resistance of the third resistor.


1
Expert's answer
2022-01-28T12:51:00-0500

Given:

"V=100\:\rm V"

"P_1=1600\:\rm W"

"P_2=1200\:\rm W"

"t=\rm 10\: min=600\: s"

"W=276*10^4\: \rm J"


The net power absorbed from supply

"P=P_1+P_2+P_3=W\/t"

The power absorbed by third resistor

"P_3=W\/t-P_1+P_2=V^2\/R_3"

Hence, the resistance of third resistor

"R_3=\frac{V^2}{W/t-P_1-P_2}"

"R_3=\frac{100^2}{276*10^4/600-1600-1200}=5.6\:\rm \Omega"